# From Language to Cognition: How LLMs Outgrow the Human Language Network

### v1

@@ Abstract
Large language models (LLMs) exhibit remarkable similarity to neural activity in the human language network. However, the key properties of language shaping brain-like representations, and their evolution during training as a function of different tasks remain unclear. We here benchmark 34 training checkpoints spanning 300B tokens across 8 different model sizes to analyze how brain alignment relates to linguistic competence. Specifically, we find that brain alignment tracks the development of formal linguistic competence—i.e., knowledge of linguistic rules—more closely than functional linguistic competence. While functional competence, which involves world knowledge and reasoning, continues to develop throughout training, its relationship with brain alignment is weaker, suggesting that the human language network primarily encodes formal linguistic structure rather than broader cognitive functions. We further show that model size is not a reliable predictor of brain alignment when controlling for feature size and find that the correlation between next-word prediction, behavioral alignment and brain alignment fades once models surpass human language proficiency. Finally, using the largest set of rigorous neural language benchmarks to date, we show that language brain alignment benchmarks remain unsaturated, highlighting opportunities for improving future models. Taken together, our findings suggest that the human language network is best modeled by formal, rather than functional, aspects of language.
Keywords: Language; Human Language Network; LLMs; Brain Alignment; Behavioral Alignment
@@ Introduction
Figure 1: Model Alignment with the Human Language Network is Primarily Driven by Formal Over Functional Linguistic Competence. (a) Average brain alignment across five Pythia models and five brain recording datasets, normalized by cross-subject consistency, throughout training. (b) Average normalized accuracy of the same models on formal linguistic competence benchmarks (two benchmarks). (c) Average normalized accuracy on functional linguistic competence benchmarks (six benchmarks). The results show that models first excel at formal linguistic tasks (such as judging the grammaticality of a sentence), aligning with the human language network, before later improving on functional benchmarks (such as reasoning about the world) but without further improvements in brain alignment. The x-axis is logarithmically spaced up to 16B tokens, capturing early training dynamics, and then evenly spaced every 20B tokens from 20B to  300B tokens. The vertical black line is at 16B tokens.
@@
Deciphering the brain’s algorithms underlying our ability to process language and communicate is a core goal in neuroscience. Human language processing is supported by the brain’s language network (LN), a set of left-lateralized fronto-temporal regions in the brain (Binder et al., 1997; Bates et al., 2003; Gorno‐Tempini et al., 2004; Price, 2010; Fedorenko, 2014; Hagoort, 2019) that respond robustly and selectively to linguistic input (Fedorenko et al., 2024). Driven by recent advances in machine learning, large language models (LLMs) trained via next-word prediction on large corpora of text are now a particularly promising model family to capture the internal processes of the LN. In particular, when these models are exposed to the same linguistic stimuli (e.g., sentences or narratives) as human participants during neuroimaging and electrophysiology experiments, they account for a substantial portion of neural response variance (Schrimpf et al., 2021; Caucheteux and King, 2022; Goldstein et al., 2022; Pasquiou et al., 2022; Aw et al., 2023; Tuckute et al., 2024a; AlKhamissi et al., 2024; Rathi et al., 2025).
Dataset Modality Presentation Context Stimulus Example
Pereira2018 fMRI Reading Short Accordions produce sound with bellows …
Blank2014 fMRI Listening Long A clear and joyous day it was and out on the wide …
Fedorenko2016 ECoG Reading Short ‘ALEX’, ‘WAS’, ‘TIRED’, ‘SO’, ‘HE’, ‘TOOK’, …
Tuckute2024 fMRI Reading Short The judge spoke, breaking the silence.
Narratives fMRI Listening Long Okay so getting back to our story about uh Lucy …
Futrell2018 Reading Times Reading Long A clear and joyous day it was and out on the wide …
Table 1: Datasets Used for Evaluating Model Alignment. Neuroimaging datasets were collected via either functional magnetic resonance imaging (fMRI) or electrocorticography (ECoG). Text stimuli range from short sentences (Fedorenko2016, Tuckute2024) to paragraphs (Pereira2018) and entire stories (Blank2014, Narratives, Futrell2018) and were presented either visually or auditorily. Futrell2018 is a behavioral dataset.
Key Questions and Contributions
@@
This work investigates four key questions, all aimed at distilling why LLM aligns to brain responses. Specifically, we investigate the full model development cycle as a combination of model architecture (structural priors) and how linguistic competence emerges across training (developmental experience). We ask: (1) What drives brain alignment in untrained models? (2) Is brain alignment primarily linked to formal or functional linguistic competence (Mahowald et al., 2024)? (3) Does model size or next-word prediction explain brain alignment? (4) Do current LLMs fully account for the explained variance in brain alignment benchmarks? To answer these questions, we introduce a rigorous brain-scoring framework to conduct a controlled and large-scale analysis of LLM brain alignment. Our findings reveal that the initial brain alignment of models with untrained parameters is driven by context integration in sequence-based models. During training, alignment primarily correlates with formal linguistic competence, which saturates early on in training ($\sim 4$B tokens), consistent with a plateauing of model-to-brain alignment. Functional competence, on the other hand, emerges later in training, tracks brain alignment less strongly, and continues to grow even when brain alignment has already saturated. This disconnect later in training is further exemplified by a fading of the correlation between models’ brain alignment and their next-word-prediction performance, as well as their behavioral alignment. Further, model size is not a reliable predictor of brain alignment when controlling for the number of features, challenging the assumption that larger models necessarily resemble the brain more. Finally, we show that current brain alignment benchmarks remain unsaturated, indicating that LLMs can still be improved in their ability to model human language processing.
Preliminaries & Related Work
@@ A Primer on Language in the Human Brain
The human language network (LN) is a set of left-lateralized frontal and temporal brain regions supporting language. These regions are functionally defined by contrasting responses to language inputs over perceptually matched controls (e.g., lists of non-words) (Fedorenko et al., 2010). The language network exhibits remarkable selectivity for language processing compared to various non-linguistic inputs and tasks, such as music perception (Fedorenko et al., 2012; Chen et al., 2023) or arithmetic computation (Fedorenko et al., 2011; Monti et al., 2012) (for review, see (Fedorenko et al., 2024)) and the language network only shows weak responses when participants comprehend or articulate meaningless non-words (Fedorenko et al., 2010; Hu et al., 2023). This selectivity profile is supported by extensive neuroimaging research and further corroborated by behavioral evidence from aphasia studies: when brain damage is confined to language areas, individuals lose their linguistic abilities while retaining other skills, such as mathematics (Benn et al., 2013; Varley et al., 2005), general reasoning (Varley and Siegal, 2000), and theory of mind (Siegal and Varley, 2006).
@@
Prior work has shown that the internal representations of certain artificial neural networks resemble those in the brain. This alignment was initially observed in the domain of vision (Yamins et al., 2014; Khaligh-Razavi and Kriegeskorte, 2014; Cichy et al., 2016; Schrimpf et al., 2018, 2020; Cadena et al., 2019; Kubilius et al., 2019; Nayebi et al., 2018; Kietzmann et al., 2019; Zhuang et al., 2021) and has more recently been extended to auditory processing (Kell et al., 2018; Tuckute et al., 2023; Koumura et al., 2023) and language processing (Schrimpf et al., 2021; Caucheteux and King, 2022; Goldstein et al., 2022; Kauf et al., 2023; Hosseini et al., 2024; Aw et al., 2023; AlKhamissi et al., 2024; Tuckute et al., 2024b; Rathi et al., 2025).
@@
Recent work in vision neuroscience has shown that untrained convolutional networks can yield high brain alignment to recordings in the visual ventral stream without the need for experience-dependent training (Geiger et al., 2022; Kazemian et al., 2024). Other works have investigated the inductive biases in different architectures and initializations in models of visual processing (Cichy et al., 2016; Cadena et al., 2019; Geiger et al., 2022), speech perception (Millet and King, 2021; Tuckute et al., 2023), and language (Schrimpf et al., 2021; Pasquiou et al., 2022; Hosseini et al., 2024), highlighting that randomly initialized networks are not random functions (Teney et al., 2024).
Methods
Benchmarks for Brain Alignment
@@ Methods
The neuroimaging datasets used in this work can be categorized along three dimensions: the imaging modality, the context length, and the modality through which the language stimulus was presented to human participants (auditory or visual). Table 1 provides an overview of all datasets in this study. We only consider neural units (electrodes, voxels, or regions) associated with the brain’s language network that were localized by their respective authors using the method described in the Functional Localization section, as packaged in Brain-Score Schrimpf et al. (2020, 2021). However, for the Narratives dataset, functional localization was not performed—to approximate the language regions, we leverage a probabilistic atlas for the human language system (Lipkin et al., 2022). Following the functional localization approach, we extract the top-10% language-selective voxels (from the probabilistic atlas) within anatomically defined language parcels. In an additional analysis, we investigate alignment with language behavior using the Futrell et al. (2018) dataset, which consists of self-paced per-word human reading times. See Appendix A for details of each dataset. To the best of our knowledge, this study examines the largest number of benchmarks compared to previous work, providing a more comprehensive and reliable foundation for understanding what drives brain alignment in LLMs. By analyzing diverse datasets, we ensure that our conclusions about brain-aligned models are not limited to specific stimuli or experimental conditions, leading to more generalizable insights.
@@
Following standard practice in measuring brain alignment, we train a ridge regression model to predict brain activity from model representations, using the same linguistic stimuli presented to human participants in neuroimaging studies (Schrimpf et al., 2020, 2021). We then measure the Pearson correlation between the predicted brain activations and the actual brain activations of human participants on a held-out set. This process is repeated over $k$ cross-validation splits, and we report the average (mean) Pearson correlation as our final result. We refer to this metric as Linear Predictivity. In the following section, we demonstrate why other metrics such as Centered Kernel Alignment (CKA; Kornblith et al., 2019) and Representational Similarity Analysis (RSA; Kriegeskorte et al., 2008) are not suitable measures for brain alignment on current language datasets.
Figure 2: Evaluating Brain Alignment with Linear Predictivity and No Contextualization is Most Stringent. (a) Average brain alignment across 8 Pythia models under three conditions: (1) a pretrained model processing the original stimuli, (2) a pretrained model processing random sequences of the same length (averaged over five random seeds) as a control condition, and (3) the model with untrained parameters processing the original stimuli. The linear predictivity metric differentiates between meaningful and random stimuli most strongly, while RSA and CKA overestimate alignment. (b) Brain alignment on the Pereira2018 dataset under two cross-validation schemes: with contextualization (random sentence split) and without contextualization (story-based split). Allowing contextualization makes it easier for untrained models to predict brain data, while enforcing stronger generalization via a story-based split reveals a significant drop in untrained model alignment.
@@
To assess the reliability of our datasets and account for the inherent noise in brain recordings, we compute a cross-subject consistency score (Feather et al., 2025), also referred to as the noise ceiling (Schrimpf et al., 2021). The consistency score is estimated by predicting the brain activity of a held-out subject using data from all other subjects, through 10-fold cross-validation of all subjects. To obtain a conservative ceiling estimate, we extrapolate subject pool sizes and report the final value based on extrapolation to infinitely many subjects. For Tuckute2024 we use the theoretical estimate provided by (Tuckute et al., 2024b). The consistency scores are provided in the Appendix. To aggregate scores across benchmarks, we normalize each model’s Pearson correlation ($r$) score for Linear Predictivity by the cross-subject consistency estimate, using the formula: (normalized score=raw $\textnormal{normalized score}=\frac{\textnormal{raw score}}{\textnormal{consistency}}$). The final brain alignment score for each model is reported as the average across all benchmarks. Otherwise, when reporting raw brain alignment, we compute the mean Pearson correlation across datasets without normalization.
Benchmarks for Formal & Functional Competence
There is substantial evidence in neuroscience research that formal and functional linguistic competence are governed by distinct neural mechanisms Mahowald et al. (2024). Formal linguistic competence pertains to the knowledge of linguistic rules and patterns, while functional linguistic competence involves using language to interpret and interact with the world. Therefore, to accurately track the evolution of each type of competence during training, we here focus on benchmarks that specifically target these cognitive mechanisms in LLMs.
@@
We evaluate all benchmarks in a zero-shot setting, using surprisal as the evaluation metric—where the model’s prediction is determined by selecting the most probable candidate, as packaged in language model evaluation harness (Gao et al., 2024). We report accuracy normalized by chance performance, where 0% indicates performance at the random chance level.
@@
Large Language Models (LLMs)
Throughout this work, we use eight models from the Pythia model suite (Biderman et al., 2023), spanning a range of sizes: {14M, 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B}. Each model is evaluated across 34 training checkpoints, spanning approximately 300B tokens. These checkpoints include the untrained model, the final trained model, and 16 intermediate checkpoints that are logarithmically spaced up to 128B tokens. The remaining 14 checkpoints are evenly spaced every 20B tokens from 20B to 280B tokens, ensuring a comprehensive analysis of alignment trends throughout training. Since smaller models fail to surpass chance performance on many functional benchmarks, we exclude 14M, 70M, 160M from analyses that compare brain alignment with functional benchmark performance.
Rigorous brain-scoring
While substantial progress has been made in measuring alignment between LLM representations and neural activity, there remains no standardized approach for comparing results across different datasets and experimental conditions. Here, we aim to establish a set of desiderata for evaluating brain alignment. For a model to be considered truly brain-aligned, two key criteria must be met. First, high alignment scores should indicate that the model captures stimulus-driven responses—meaning that when presented with a random sequence of tokens, alignment should drop significantly compared to original linguistic stimuli. Second, a brain-aligned model should generalize effectively to new linguistic contexts rather than overfitting to specific examples. Below, we address these two points to justify our choice of metric and cross-validation scheme for each dataset. For all benchmarks, we localize language-selective units, which is consistent with neural site selection in neuroscience experiments and allows for fair comparisons across models irrespective of model size.
Figure 3: Context Integration drives Brain Alignment of Untrained Models. (a) Brain alignment of models with untrained parameters across five random initializations, showing that sequence-based models (GRU, LSTM, Transformers, and mean pooling) achieve higher alignment than models that rely solely on the last token representation (Linear, MLP), highlighting the importance of temporal integration. (b) Ablation study of architectural components in a single untrained Transformer-v2 block, demonstrating that attention mechanisms combined with positional encoding yield the highest brain alignment. (c) Diagram of the Transformer block architecture used in (b), with components grouped into attention (lower box) and MLP (upper box). (d) The average performance of five Pythia models with untrained parameters on formal and functional linguistic competence benchmarks, showing that formal competence exceeds chance level even in untrained parameter models.
Functional Localization
The human language network (LN) is defined functionally which means that units are chosen according to a ‘localizer’ experiment (Saxe et al., 2006). Specifically, the LN is the set of neural units (e.g., voxels/electrodes) that are more selective to sentences over a perceptually-matched control condition (Fedorenko et al., 2010). When selecting units from artificial models for comparison against LN units, previous work selected output units from an entire Transformer block based on brain alignment scores (Schrimpf et al., 2021). However, LLMs learn diverse concepts and behaviors during their considerable pretraining, not all of which are necessarily related to language processing, e.g., storage of knowledge (AlKhamissi et al., 2022) and the ability to perform complex reasoning (Huang and Chang, 2023). Therefore, we here follow the method proposed by AlKhamissi et al. (2024) that identifies language units in LLMs using functional localization as is already standard in neuroscience. This approach offers a key advantage: it enables direct comparisons across models by selecting a fixed set of units, identified through the independent localizer experiment. In this work, we localize $128$ units for all models unless otherwise specified, and we show in the Appendix that the results hold when selecting a different number of units.
Robust Metrics and Generalization Tests

### v2

@@ Abstract
Large language models (LLMs) exhibit remarkable similarity to neural activity in the human language network. However, the key properties of language underlying this alignment—and how brain-like representations emerge and change across training—remain unclear. We here benchmark 34 training checkpoints spanning 300B tokens across 8 different model sizes to analyze how brain alignment relates to linguistic competence. Specifically, we find that brain alignment tracks the development of formal linguistic competence—i.e., knowledge of linguistic rules—more closely than functional linguistic competence. While functional competence, which involves world knowledge and reasoning, continues to develop throughout training, its relationship with brain alignment is weaker, suggesting that the human language network primarily encodes formal linguistic structure rather than broader cognitive functions. Notably, we find that the correlation between next-word prediction, behavioral alignment, and brain alignment fades once models surpass human language proficiency. We further show that model size is not a reliable predictor of brain alignment when controlling for the number of features. Finally, using the largest set of rigorous neural language benchmarks to date, we show that language brain alignment benchmarks remain unsaturated, highlighting opportunities for improving future models. Taken together, our findings suggest that the human language network is best modeled by formal, rather than functional, aspects of language. [^1] Project Page: language-to-cognition.epfl.ch
From Language to Cognition:
How LLMs Outgrow the Human Language Network
Badr AlKhamissi1   Greta Tuckute2   Yingtian Tang1   Taha Binhuraib3 Antoine Bosselut∗,1Martin Schrimpf∗,1 1EPFL  2MIT  3Georgia Institute of Technology
@@ 1 Introduction
Figure 1: Model Alignment with the Human Language Network is Primarily Driven by Formal than Functional Linguistic Competence. (a) Average brain alignment across five Pythia models and five brain recording datasets, normalized by cross-subject consistency, throughout training. (b) Average normalized accuracy of the same models on formal linguistic competence benchmarks (two benchmarks). (c) Average normalized accuracy on functional linguistic competence benchmarks (six benchmarks). The x-axis is logarithmically spaced up to 16B tokens, capturing early training dynamics, and then evenly spaced every 20B tokens from 20B to ~300B tokens.
@@
Deciphering the brain’s algorithms underlying our ability to process language and communicate is a core goal in neuroscience. Human language processing is supported by the brain’s language network (LN), a set of left-lateralized fronto-temporal regions in the brain (Binder et al., 1997; Bates et al., 2003; Gorno‐Tempini et al., 2004; Price, 2010; Fedorenko, 2014; Hagoort, 2019) that respond robustly and selectively to linguistic input (Fedorenko et al., 2024a). Driven by recent advances in machine learning, large language models (LLMs) trained via next-word prediction on large corpora of text are now a particularly promising model family to capture the internal processes of the LN. In particular, when these models are exposed to the same linguistic stimuli (e.g., sentences or narratives) as human participants during neuroimaging and electrophysiology experiments, they account for a substantial portion of neural response variance (Schrimpf et al., 2021; Caucheteux and King, 2022; Goldstein et al., 2022; Pasquiou et al., 2022; Aw et al., 2023; Tuckute et al., 2024a; AlKhamissi et al., 2025; Rathi et al., 2025).
1.1 Key Questions and Contributions
@@
This work investigates four key questions, all aimed at distilling why LLM aligns to brain responses. Specifically, we investigate the full model development cycle as a combination of model architecture (structural priors) and how linguistic competence emerges across training (developmental experience). We ask: (1) What drives brain alignment in untrained models? (2) Is brain alignment primarily linked to formal or functional linguistic competence (Mahowald et al., 2024)? (3) Do language models diverge from humans as they surpass human-level prediction? (4) Do current LLMs fully account for the explained variance in brain alignment benchmarks? To answer these questions, we introduce a rigorous brain-scoring framework to conduct a controlled and large-scale analysis of LLM brain alignment.
Our findings reveal that the initial brain alignment of models with untrained parameters is driven by context integration. During training, alignment primarily correlates with formal linguistic competence—tasks that probe mastery of grammar, syntax, and compositional rules, such as identifying subject–verb agreement, parsing nested syntactic structures, or completing well-formed sentences. This competence saturates relatively early in training ($\sim 4$B tokens), consistent with a plateauing of model-to-brain alignment. Functional linguistic competence, in contrast, concerns how language is used in context to convey meaning, intent, and social/pragmatic content—for example, tasks involving discourse coherence, reference resolution, inference about speaker meaning, or interpreting figurative language. Functional competence emerges later in training, tracks brain alignment less strongly, and continues to grow even after alignment with the language network has saturated.
This disconnect later in training is further exemplified by a fading of the correlation between models’ brain alignment and their next-word-prediction performance, as well as their behavioral alignment. Further, we show that model size is not a reliable predictor of brain alignment when controlling for the number of features, challenging the assumption that larger models necessarily resemble the brain more. Finally, we demonstrate that current brain alignment benchmarks remain unsaturated, indicating that LLMs can still be improved to model human language processing.
2 Preliminaries & Related Work
@@ A Primer on Language in the Human Brain
The human language network (LN) is a set of left-lateralized frontal and temporal brain regions supporting language. These regions are functionally defined by contrasting responses to language inputs over perceptually matched controls (e.g., lists of non-words) (Fedorenko et al., 2010). The language network exhibits remarkable selectivity for language processing compared to various non-linguistic inputs and tasks, such as music perception (Fedorenko et al., 2012; Chen et al., 2023) or arithmetic computation (Fedorenko et al., 2011; Monti et al., 2012) (for review, see Fedorenko et al. (2024a)) and the language network only shows weak responses when participants comprehend or articulate meaningless non-words (Fedorenko et al., 2010; Hu et al., 2023). This selectivity profile is supported by extensive neuroimaging research and further corroborated by behavioral evidence from aphasia studies: when brain damage is confined to language areas, individuals lose their linguistic abilities while retaining other skills, such as mathematics (Benn et al., 2013; Varley et al., 2005), general reasoning (Varley and Siegal, 2000), and theory of mind (Siegal and Varley, 2006).
@@
Prior work has shown that the internal representations of certain artificial neural networks resemble those in the brain. This alignment was initially observed in the domain of vision (Yamins et al., 2014; Khaligh-Razavi and Kriegeskorte, 2014; Cichy et al., 2016; Schrimpf et al., 2018, 2020; Cadena et al., 2019; Kubilius et al., 2019; Zhuang et al., 2021) and has more recently been extended to auditory processing (Kell et al., 2018; Tuckute et al., 2023; Koumura et al., 2023) and language processing (Schrimpf et al., 2021; Caucheteux and King, 2022; Goldstein et al., 2022; Kauf et al., 2023; Hosseini et al., 2024; Aw et al., 2023; AlKhamissi et al., 2025; Tuckute et al., 2024b; Rathi et al., 2025).
@@
Recent work in vision neuroscience has shown that untrained convolutional networks can yield high brain alignment to recordings in the visual ventral stream without the need for training (Geiger et al., 2022; Kazemian et al., 2024). Other works have investigated the inductive biases in different architectures and initializations in models of visual processing (Cichy et al., 2016; Cadena et al., 2019; Geiger et al., 2022), speech perception (Millet and King, 2021; Tuckute et al., 2023), and language (Schrimpf et al., 2021; Pasquiou et al., 2022; Hosseini et al., 2024), highlighting that randomly initialized networks are not random functions (Teney et al., 2024).
3 Methods
3.1 Benchmarks for Brain Alignment
@@
The neuroimaging datasets used in this work can be categorized along three dimensions: the imaging modality, the context length of the experimental materials, and the modality through which the language stimulus was presented to human participants (auditory or visual). Table 1 in Appendix A provides an overview of all datasets in this study. To focus specifically on language, we consider neural units (electrodes, voxels, or regions) associated with the brain’s language network, as localized by the original dataset authors using the method described in the Section 3.2 and implemented in Brain-Score Schrimpf et al. (2020, 2021) (however, see Appendix J for control brain regions). An exception is the Narratives dataset, which lacks functional localization. We here approximate the language regions using a probabilistic atlas of the human language network (Lipkin et al., 2022), extracting the top-10% language-selective voxels (from the probabilistic atlas) within anatomically defined language parcels, in line with the functional localization procedure used in the other datasets. In an additional analysis, we investigate model alignment with language behavior using the Futrell et al. (2018) dataset, which contains self-paced, per-word human reading times. See Appendix A for details of each dataset. To the best of our knowledge, this study examines the largest number of benchmarks compared to previous work, providing a more comprehensive and reliable foundation for identifying the properties that drive brain alignment in LLMs. The diversity of datasets ensures that our conclusions generalize beyond specific experimental stimuli and paradigms.
@@
Following standard practice in measuring brain alignment, we train a ridge regression model to predict brain activity from model representations, using the same linguistic stimuli presented to human participants in neuroimaging studies (Schrimpf et al., 2020, 2021). We then measure the Pearson correlation between the predicted brain activations and the actual brain activations of human participants on a held-out set that covers entirely different stories or topics (see Section 4). This process is repeated over $k$ cross-validation splits, and we report the average (mean) Pearson correlation as our final result. We refer to this metric as Linear Predictivity. In Section 5.1, we demonstrate why other metrics such as Centered Kernel Alignment (CKA; Kornblith et al., 2019) and Representational Similarity Analysis (RSA; Kriegeskorte et al., 2008) are not suitable measures for brain alignment on current language datasets.
@@
To assess the reliability of our datasets and account for the inherent noise in brain recordings, we compute a cross-subject consistency score (Feather et al., 2025), also referred to as the noise ceiling (Schrimpf et al., 2021). The consistency score is estimated by predicting the brain activity of a held-out subject using data from all other subjects, through 10-fold cross-validation of all subjects. To obtain a conservative ceiling estimate, we extrapolate subject pool sizes and report the final value based on extrapolation to infinitely many subjects. For Tuckute2024 we use the theoretical estimate provided by (Tuckute et al., 2024b). Consistency scores are provided in Appendix K. To aggregate scores across benchmarks, we normalize each model’s Pearson correlation ($r$) score for Linear Predictivity by the cross-subject consistency estimate, using the formula: ($\textnormal{normalized score}=\frac{\textnormal{raw score}}{\textnormal{consistency}}$). The final alignment score for each model is reported as the average across all benchmarks. Otherwise, when reporting raw alignment, we compute the mean Pearson correlation across datasets without normalization.
3.2 Functional Localization
The human language network (LN) is defined functionally which means that units are chosen according to a ‘localizer’ experiment (Saxe et al., 2006). Specifically, the LN is the set of neural units (e.g., voxels/electrodes) that are more selective to sentences over a perceptually-matched control condition (Fedorenko et al., 2010). When selecting units from artificial models for comparison against LN units, previous work selected output units from an entire Transformer block based on brain alignment scores (Schrimpf et al., 2021). However, LLMs learn diverse concepts and behaviors during their considerable pretraining, not all of which are necessarily related to language processing, e.g., storage of knowledge (AlKhamissi et al., 2022) and the ability to perform complex reasoning (Huang and Chang, 2023). Therefore, we here follow the method proposed by AlKhamissi et al. (2025) that identifies language units in LLMs using functional localization as is already standard in neuroscience. This approach offers a key advantage: it enables direct comparisons across models by selecting a fixed set of units, identified through the independent localizer experiment. In this work, we localize $128$ units for all models unless otherwise specified, and we show in Appendix H that the results hold when selecting a different number of units.
Figure 2: Context Integration drives Brain Alignment of Untrained Models. (a) Sequence-based models (GRU, LSTM, Transformers, and mean pooling) achieve higher brain alignment than models that rely solely on the last token representation (Linear, MLP), highlighting the importance of temporal integration. Error bars report five random initializations in all subplots. (b) Ablation study of architectural components in a single untrained Transformer-v2 block, demonstrating that attention mechanisms combined with positional encoding yield the highest brain alignment. (c) Diagram of the Transformer block architecture used in (b), with components grouped into attention (lower box) and MLP (upper box). (d) The average performance of five Pythia models with untrained parameters on formal and functional linguistic competence benchmarks, showing that formal competence exceeds chance level even in untrained parameter models.
3.3 Benchmarks for Linguistic Competence
There is substantial evidence in neuroscience research that formal and functional linguistic competence are governed by distinct neural mechanisms Mahowald et al. (2024); Fedorenko et al. (2024a, b). Formal linguistic competence pertains to the knowledge of linguistic rules and patterns, while functional linguistic competence involves using language to interpret and interact with the world. Therefore, to accurately track the evolution of each type of competence during training, we focus on benchmarks that specifically target these cognitive capacities in LLMs.
@@ 3.3 Benchmarks for Linguistic Competence
Inline with prior work, we evaluate all benchmarks in a zero-shot setting, using surprisal as the evaluation metric. where the model’s prediction is determined by selecting the most probable candidate, as packaged in the language model evaluation harness (Gao et al., 2024). We report accuracy normalized by chance performance, where 0% indicates performance at the random chance level.
@@
3.4 Large Language Models (LLMs)
Throughout this work, we use eight models from the Pythia model suite (Biderman et al., 2023), spanning a range of sizes: {14M, 70M, 160M, 410M, 1B, 1.4B, 2.8B, 6.9B}. Each model is evaluated across 34 training checkpoints, spanning approximately 300B tokens. These checkpoints include the untrained model, the final trained model, and 16 intermediate checkpoints that are logarithmically spaced up to 128B tokens. The remaining 14 checkpoints are evenly spaced every 20B tokens from 20B to 280B tokens, ensuring a comprehensive analysis of alignment trends throughout training. Since smaller models fail to surpass chance performance on many functional benchmarks, we exclude 14M, 70M, 160M from analyses that compare brain alignment with functional performance.
4 Rigorous Brain-Scoring
While substantial progress has been made in measuring alignment between LLM representations and neural activity, there’s no standard for comparing brain alignment across datasets and conditions. Therefore, to ensure we perform meaningful inferences, we propose two criteria: (1) alignment should reflect stimulus-driven responses, dropping for random token sequences; and (2) models should generalize to new linguistic contexts. We justify our metrics and cross-validation choices accordingly. For all benchmarks, we identify language-selective units to ensure fair model comparisons, consistent with neural site selection in neuroscience AlKhamissi et al. (2025).
4.1 Robust Metrics and Generalization Tests